# A Radiance Field Loss for Fast and Simple Emissive Surface Reconstruction

### v1

@@ Abstract.
We present a fast and simple technique to convert images into an emissive surface-based scene representation. Building on existing emissive volume reconstruction algorithms, we introduce a subtle yet impactful modification of the loss function requiring changes to only a few lines of code: instead of integrating the radiance field along rays and supervising the resulting images, we project the training images into the scene to directly supervise the spatio-directional radiance field.
The primary outcome of this change is the complete removal of alpha blending and ray marching from the image formation model, instead moving these steps into the loss computation. In addition to promoting convergence to surfaces, this formulation assigns explicit semantic meaning to 2D subsets of the radiance field, turning them into well-defined emissive surfaces. We finally extract a level set from this representation, which results in a high-quality emissive surface model.
@@
copyright: nonejournalyear: 2025doi: XXXXXXX.XXXXXXXconference: ; ; journal: TOG
@@ 1. Introduction
The resulting radiance field loss considers each point along a ray as a surface candidate, individually optimized to match that ray’s pixel color, leading to the desired distribution over surfaces. One benefit is that points along a ray receive independent gradients, allowing the color or density to simultaneously increase at a point and decrease at another. This is notably different to the volumetric approach, which integrates the color along the ray prior to the loss computation (see Figure 1, left). That is, with volume reconstruction, all points along a ray receive a gradient with the same sign if their integrated color is too dark or bright, leading to correlated adjustments.
Interestingly, our proposed radiance field loss gives rise to equations remarkably similar to those of volumetric reconstruction methods (see Figure 2). In practical terms, this means that our method is simple to integrate into existing volumetric frameworks. It also means that we inherit many advantages of these prior works, without having to resort to additional heuristics to extract a surface. While we have not focused on competing with existing methods in terms of metrics, our proof of concept implementation in Instant NGP [Müller et al., 2022] consists of a few modified lines of code in the core algorithm and runs at roughly the same speed (in terms of PSNR vs. time) while producing surfaces whose PSNR is, on average, only $0.1$ dB lower than that of the volumetric approach.
@@ 2. Related work
The first works on differentiable rendering embraced the high-level approach of optimizing an initial guess of a shape via gradient descent [Loper and Black, 2014], variously representing the surface using SDF level sets [Zhang et al., 2021; Vicini et al., 2022], triangle meshes [Nicolet et al., 2021], points [Chen et al., 2024b], or hybrids [Munkberg et al., 2022]. Regardless of the underlying representation, it remains challenging to achieve satisfactory results in this way: this is partly due to the complex loss landscape of an evolving surface, and partly due to the numerical difficulties of computing visibility-induced gradients [Loubet et al., 2019; Zhang et al., 2020, 2023]. Without intricate special-case handling, the optimization often fails when topological changes are required [Mehta et al., 2023], or when the surface does not overlap with the target shape [Xing et al., 2023]. Our work sidesteps these limitations by replacing the surface boundary with a distribution over surfaces.
@@
After the advent of emissive volume reconstruction (NeRF) for novel view synthesis [Mildenhall et al., 2020], researchers developed various regularizers and parameterizations of emissive volumes to ensure that their level sets yield plausible geometry [Wang et al., 2021; Yariv et al., 2021, 2023]. Surfaces can then be extracted using established algorithms like marching cubes. However, while efficient NeRF implementations reconstruct in seconds to minutes [Müller et al., 2022], methods in the aforementioned line of work require hours of computation [Li et al., 2023] or result in substantially reduced quality [Wang et al., 2023]. In contrast, our method largely preserves the reconstruction speed and quality of the baseline NeRF method.
@@
Several prior works conceptualized volumetric reconstruction as optimizing a distribution over surfaces [Seyb et al., 2024; Miller et al., 2024; Wang et al., 2021]. These methods however represent objects as the union of multiple interacting surfaces (whose contributions are integrated along the ray), which conflicts with our end goal of extracting a single surface to model an object geometry. Instead, we build upon the “many worlds” concept proposed by Zhang et al. [2024], which considers a distribution of non-interacting surfaces, and apply it to the problem of emissive surface reconstruction. We show how, in this context, the many worlds concept gives rise to a simple equation dual to the one used in NeRF frameworks; see Figure 2.
@@ 4.1. Novel view synthesis
Table 1 shows that our method achieves a PSNR only slightly lower than NeRF when trained on the MipNeRF360 dataset with the same hyperparameters, despite employing a surface-based representation. A PSNR gap is expected, as volume representations offer inherently more degrees of freedom that can be repurposed to handle reflections and fine details. This gap is larger in outdoor scenes, in which the far-field is especially hard to model with surfaces.
When training the relaxed variant that switches to volumetric rendering in hard regions in Instant NGP, the visual quality slightly surpasses the NeRF baseline (we measured indoor and outdoor mean PSNRs of $22.62$ dB and $29.41$ dB, respectively). This is because our method promotes surface-like, sparse distributions, leading to more empty space that the underlying rendering algorithm can skip. Consequently, at equal batch size, Instant NGP automatically spawns more rays when using our method, thereby covering more reference pixels per batch, in turn leading to a better reconstruction. When the ray count is restricted to match NeRF, the relaxed variant delivers PSNR results that are approximately equal.
@@
Our implementation builds on the Instant NGP codebase, which ray-marches fields ($\alpha_{\mathbf{p}},L_{\mathbf{p}}$) represented using an interpolated hash grid lookup combined with a lightweight MLP. We repurpose this ray-marching code for surface rendering by returning the color of the first sample with an occupancy value exceeding $0.5$. This straightforward modification results in an 2.4×2.4$\times$2.4 × average speedup in frames per second (FPS) across MipNeRF360 scenes compared to the baseline. An average speedup of 2.0×2.0$\times$2.0 × is achieved for the relaxed version of our method, as most of the scene remains surface-like.
@@
Table 1. PSNR comparison. We integrate our loss into Instant NGP and ZipNeRF and train on the MipNeRF360 dataset using default hyperparams.
Instant NGP codebase ZipNeRF codebase
Ours NeRF Ours NeRF
Indoor mean 22.41 dB 22.47 dB 24.06 dB 25.24 dB
Outdoor mean 29.02 dB 29.19 dB 29.73 dB 31.45 dB
@@ 4.2. Geometry reconstruction
Laplacian regularization is only the first step in a long sequence of potential improvements that can further boost geometric reconstruction quality. These include multi-view consistency losses [Fu et al., 2022; Chen et al., 2024a] to reduce ambiguities in regions with limited observations, or the TSDF algorithm [Izadi et al., 2011] that helps extract smooth meshes while removing unnecessary geometry. We do not intend to compete on geometric reconstruction metrics and have not incorporated such extensions, which would detract from the simplicity of the presented idea.
Figure 7. Straightforward extraction. Since our algorithm does not use an intermediate volume representation, efficient surface extraction is possible at any point. At equal time, a fast NeuS baseline [Wang et al., 2023] still models the scene as a fuzzy volume, and a surface cannot be confidently extracted.
@@ 5.2. Relation to many-worlds inverse rendering
Our method builds on the core idea of Zhang et al. [2024] (we refer to their method as PBR-MW), namely that surface distributions can be optimized more directly without involving exponential volumes. We reconstruct purely emissive objects, while PBR-MW handles differentiable shadowing and interreflection to reconstruct reflecting objects in scenes with global illumination. Viewed superficially, our method could be mistaken for a stripped down version of PBR-MW.
Our contribution lies in leveraging this simplicity to develop a specialized method. We identify and implement optimizations unique to emissive surfaces to fully realize the potential of the many-worlds idea.
In emissive surface rendering, image formation is a direct 1:1 mapping between a ray and the nearest intersected emissive surface, while PBR-MW requires a complex nested integration over materials, lighting, and geometry. Our approach to project training images into the scene to establish a radiance field loss depends on this 1:1 mapping and does not efficiently translate to the nested integral structure of a global illumination renderer.
Another important contribution is the introduction of a stochastic background distribution, which enables topological changes and substantially improves reconstruction quality. We show how to cheaply evaluate this strategy in expectation, which is needed to maintain algorithmic parity with NeRF. The associated derivations and simplifications (Appendix A) are specific to emissive surfaces and do not transfer to PBR-MW.
@@ 5.3. Limitations and future work
As shown in Figure 8, our lightweight Laplacian regularization fails when there are insufficient observations to constrain the geometry. Using alternative regularization techniques from state-of-the-art geometry reconstruction methods could help mitigate this issue. Our method also struggles to accurately capture the appearance of conductive materials, which could be addressed by incorporating solutions from prior work [Verbin et al., 2021].
An interesting extension of our work could involve implementing a particle-based storage approach, such as Gaussian splatting [Niemeyer et al., 2020]. However, 3D Gaussians are inherently semi-transparent, which conflicts with our assumption of opacity. Future work could explore the use of opaque primitives, such as 2D disks, to replace semi-transparent particles.
@@ 6. Conclusion
The ”many worlds” paradigm—i.e., optimizing a distribution over non-interacting primitives—is relatively new in the field of differentiable rendering. In this paper, we apply it to emissive surface reconstruction, which yields a fast and simple alternative to prior works. Particularly notable is that the derivation began with an evolving surface, yet resulted in remarkably similar equations to volumetric scene reconstructions: ones where losses rather than colors are integrated along rays.
@@
Much engineering has gone into the design of optimized algorithms, regularizers, and heuristics for NeRF-based 3D reconstruction. Our hope is that a large portion of this effort will translate to the radiance field loss and yield state of the art results in the future.
@@ Acknowledgements.
Figure 9. Nature of the reconstructed occupancy. Surface rendering at varying level sets of a scene reconstructed by our method and NeRF, both implemented in Instant NGP using the same hyperparameters. (a) For our method, the surface rendering shows minimal changes across different level set thresholds, indicating that the occupancy field has converged to a near-Heaviside step function on the surface, allowing for extraction of a surface-based representation. (b) NeRF reconstructs the scene volumetrically, and any surface extracted using a level set is a poor approximation of the true color.
Figure 10. Volumetric relaxation. We compare reconstructions of our method without and with volume relaxation to NeRF, all implemented in Instant NGP using the same hyperparameters. While our method achieves comparable visual quality using a surface-based representation, we highlight a region (white arrow) where it fails to model a semi-transparent object due to the opaque surface assumption. The relaxed variant of our algorithm can recover by adopting volume rendering in such regions. Rendering the reconstructions using the same ray marching implementation leads to significant performance differences: our surface-only reconstruction is 2.6×2.6$\times$2.6 × faster than NeRF. The relaxed variant benefits from the surface representation in most regions, and is 1.7×1.7$\times$1.7 × faster.
Figure 11. Effect of Laplacian weight on geometry reconstruction. The above results demonstrate the trade-off between geometric detail and surface smoothness. For simple scenes lacking intricate features (bottom row), the reconstruction is insensitive to this hyperparameter.
Figure 12. Reconstruction showcase. Surface rendering and normals of various scenes from the DTU and BlendedMVS datasets, reconstructed with our algorithm and a decaying Laplacian. All results were generated with the same hyperparameters and a training time of $\sim 1$ minute per scene.
@@ Appendix A Derivation of our loss in NeRF form
Following the introduction of the radiance field loss and the stochastic background, a naive implementation of our method would first sample a surface $\mathcal{M}_{\text{b}}$ from a distribution as the perturbation background, and for each sampled $\mathcal{M}_{\text{b}}$, we need to sample multiple candidates to solve the non-local perturbation problem. Naively applying this strategy would result in a quadratic time complexity. In the following, we derive an expectation of losses over all potential background surfaces (Figure 13) for a specific candidate.
Let $f_{\mathbf{b}}$ be the probability distribution of the background surface along a ray, where $\int_{0}^{t_{\mathrm{max}}}f_{\mathbf{b}}(t)\,\mathrm{d}t=1$. Without loss of generality, we focus on one candidate position $\mathbf{p}$ at distance $t_{\mathbf{p}}$ along the ray. To avoid clutter, we denote the color error metric $\ell(L,~{}L_{\mathrm{target}})$ as $\ell(L)$. The expectation of all losses in the form of Equation 2, local at $\mathbf{p}$, is given by:
@@
This final result is equivalent to the one shown in Figure 2. It is now also apparent that we do not need to produce extra samples to evaluate $\mathbb{E}_{t>t_{\mathbf{p}}}[\ell(L_{t})]$, thereby reducing the algorithm’s time complexity to a linear one.
@@ Appendix B Design space of the background distribution
Section 3.3 proposes the stochastic background surface. Designing the background surface distribution $f_{\mathbf{b}}$ (Figure 13) involves a tradeoff between exploitation and exploration, offering a wide design space.
@@ Appendix C Additional experiments and results
Applying a Laplacian as a post-processing step requires numerous iterations to address these issues and may degrade geometry in other regions. In contrast, training our algorithm with an exponentially decaying Laplacian is more efficient. Consequently, the results in Figure 12 are obtained by training our algorithm with an exponentially decaying Laplacian.
@@
Figure 20 shows additional level set rendering results for an outdoor scene. Table 2 and Table 3 show the complete PSNR and SSIM results in the Instant NGP codebase. Table 4 shows the complete PSNR results in the ZipNeRF codebase.

### v2

# Radiance Surfaces: Optimizing Surface Representations with a 5D Radiance Field Loss

@@ Abstract.
We present a fast and simple technique to convert images into a radiance surface-based scene representation. Building on existing radiance volume reconstruction algorithms, we introduce a subtle yet impactful modification of the loss function requiring changes to only a few lines of code: instead of integrating the radiance field along rays and supervising the resulting images, we project the training images into the scene to directly supervise the spatio-directional radiance field.
The primary outcome of this change is the complete removal of alpha blending and ray marching from the image formation model, instead moving these steps into the loss computation. In addition to promoting convergence to surfaces, this formulation assigns explicit semantic meaning to 2D subsets of the radiance field, turning them into well-defined radiance surfaces. We finally extract a level set from this representation, which results in a high-quality radiance surface model.
@@
copyright: acmlicensedconference: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers ; August 10–14, 2025; Vancouver, BC, Canadabooktitle: Special Interest Group on Computer Graphics and Interactive Techniques Conference Conference Papers (SIGGRAPH Conference Papers ’25), August 10–14, 2025, Vancouver, BC, Canadadoi: 10.1145/3721238.3730713isbn: 979-8-4007-1540-2/2025/08
@@ 1. Introduction
The resulting radiance field loss considers each point along a ray as a surface candidate, individually optimized to match that ray’s pixel color, leading to the desired distribution over surfaces. One benefit is that points along a ray receive independent gradients, allowing the color or density to simultaneously increase at one point and decrease at another. This is notably different from the volumetric approach, which integrates the color along the ray prior to the loss computation (see Figure 1, left). That is, with volume reconstruction, all points along a ray receive gradients with the same sign if their integrated color is too dark or bright, leading to correlated adjustments.
Interestingly, our proposed radiance field loss gives rise to equations remarkably similar to those of volumetric reconstruction methods (see Figure 2). In practical terms, this means that our method is simple to integrate into existing volumetric frameworks. It also means that we inherit many advantages of these prior works without having to resort to additional heuristics to extract a surface. While we have not focused on competing with existing methods in terms of metrics, our proof-of-concept implementation in Instant NGP [Müller et al., 2022] consists of only a few modified lines of code in the core algorithm and runs at roughly the same speed (in terms of PSNR vs. time) while producing surfaces whose PSNR is, on average, only $0.1$ dB lower than that of the volumetric baseline.
@@ 2. Related work
The first works on differentiable rendering embraced the high-level approach of optimizing an initial guess of a shape via gradient descent [Loper and Black, 2014], variously representing the surface using SDF level sets [Zhang et al., 2021; Vicini et al., 2022; Wang et al., 2024], triangle meshes [Nicolet et al., 2021], points [Chen et al., 2024b], or hybrids [Munkberg et al., 2022]. Regardless of the underlying representation, it remains challenging to achieve satisfactory results in this way: this is partly due to the complex loss landscape of an evolving surface, and partly due to the numerical difficulties of computing visibility-induced gradients [Loubet et al., 2019; Zhang et al., 2020, 2023]. Without intricate special-case handling, the optimization often fails when topological changes are required [Mehta et al., 2023], or when the surface does not overlap with the target shape [Xing et al., 2023]. Our work sidesteps these limitations by replacing the surface boundary with a distribution over surfaces.
@@
After the advent of radiance volume reconstruction (NeRF) for novel view synthesis [Mildenhall et al., 2020], researchers developed various regularizers and parameterizations of radiance volumes to ensure that their level sets yield plausible geometry [Wang et al., 2021; Yariv et al., 2021, 2023]. Surfaces can then be extracted using established algorithms like marching cubes. However, while efficient NeRF implementations reconstruct in seconds to minutes [Müller et al., 2022], methods in the aforementioned line of work require hours of computation [Li et al., 2023] or result in substantially reduced quality [Wang et al., 2023]. In contrast, our method largely preserves the reconstruction speed and quality of the baseline NeRF method.
@@
Several prior works conceptualized volumetric reconstruction as optimizing a distribution over surfaces [Seyb et al., 2024; Miller et al., 2024; Wang et al., 2021]. These methods however represent objects as the union of multiple interacting surfaces (whose contributions are integrated along the ray), which conflicts with our end goal of extracting a single surface to model an object geometry. Instead, we build upon the “many worlds” concept proposed by Zhang et al. [2024], which considers a distribution of non-interacting surfaces, and apply it to the problem of radiance surface reconstruction. We show how, in this context, the many worlds concept gives rise to a simple equation dual to the one used in NeRF frameworks; see Figure 2.
@@ 4.1. Novel view synthesis
Table 1 shows that our method achieves visual quality comparable to exponential volume reconstruction (NeRF) when trained on the MipNeRF360 dataset, using default Instant NGP hyperparameters, despite using a surface-based representation. A small PSNR gap is expected, as volume representations offer inherently more degrees of freedom that can be repurposed to model pixel-wise colors. A similar trend is observed when implementing our method in the ZipNeRF codebase, where we measured mean PSNR of $29.73$ dB for our method and $31.45$ dB for NeRF on indoor scenes, and $24.06$ dB and $25.24$ dB on outdoor scenes, respectively.
When evaluating our relaxed variant—which switches to volumetric rendering in hard regions—the visual quality slightly exceeds the NeRF baseline. This improvement arises because our method encourages surface-like, sparse distributions, resulting in more empty space that the renderer can efficiently skip. Consequently, at equal batch size, Instant NGP automatically spawns more rays when using our method, thereby covering more reference pixels per batch, in turn leading to a better reconstruction. When the ray count is restricted to match NeRF, the relaxed variant delivers results that are approximately equal.
These trends are consistent across other metrics as well. For instance, both our method and NeRF achieve SSIM scores of $0.89$ (indoor) and $0.68$ (outdoor). The full set of evaluation results is provided in the appendix.
@@
Our implementation builds on the Instant NGP codebase, which ray-marches fields ($\alpha_{\mathbf{p}},L_{\mathbf{p}}$) represented using an interpolated hash grid lookup combined with a lightweight MLP. We repurpose this ray-marching code for surface rendering by returning the color of the first sample with an occupancy value exceeding $0.5$. This straightforward modification results in a 2.4×2.4$\times$2.4 × average speedup in frames per second (FPS) across MipNeRF360 scenes compared to the baseline. An average speedup of 2.0×2.0$\times$2.0 × is achieved for the relaxed version of our method, as most of the scene remains surface-like.
@@
Table 1. Visual quality comparison. We integrate our loss into Instant NGP and train on the MipNeRF360 dataset using default hyperparams.
Indoor mean Outdoor mean
[…]
Ours 29.02 dB 0.888 $\bm{0.275}$ 22.41 dB 0.679 $\bm{0.563}$
Ours (relaxed) $\bm{29.41}\,\mathrm{\mathbf{dB}}$ $\bm{0.897}$ 0.284 $\bm{22.62}\,\mathrm{\mathbf{dB}}$ $\bm{0.690}$ 0.626
NeRF 29.19 dB 0.893 0.303 22.47 dB 0.683 0.638
@@ 4.2. Geometry reconstruction
Table 2. Average Chamfer distance comparison on the DTU dataset with NeuS [Wang et al., 2021] and NeuS2 [Wang et al., 2023].
Ours (1 min) NeuS (8 hr) NeuS2 (5 min)
CD 0.80 0.77 $\bm{0.68}$
Figure 7. Straightforward extraction. Since our algorithm does not use an intermediate volume representation, efficient surface extraction is possible at any point. At equal time, a fast NeuS2 baseline [Wang et al., 2023] still models the scene as a fuzzy volume, and a surface cannot be confidently extracted.
@@
Table 2 shows that using only minimal Laplacian regularization, our method achieves an average Chamfer distance on the DTU dataset that is just $0.12$ higher than NeuS2 [Wang et al., 2023], while reducing runtime to only 1 minute thanks to our algorithmic simplicity. The complete evaluation results are provided in the appendix. In this work, we do not intend to compete on geometric reconstruction metrics and have not incorporated other regularization extensions, which would detract from the simplicity of the presented idea. Such extensions include multi-view consistency losses [Fu et al., 2022; Chen et al., 2024a] to reduce ambiguities in regions with limited observations, or the TSDF algorithm [Izadi et al., 2011] that helps extract smooth meshes while removing unnecessary geometry.
@@ 5.2. Relation to many-worlds inverse rendering
Our method builds on the core idea of Zhang et al. [2024] (we refer to their method as PBR-MW), namely that surface distributions can be optimized more directly without involving exponential volumes. We reconstruct purely emissive objects [^2] In the equations of physically based rendering, radiance fields manifest in the emission term [Nimier-David et al., 2022]., while PBR-MW handles differentiable shadowing and interreflection to reconstruct reflecting objects in scenes with global illumination. Viewed superficially, our method could be mistaken for a stripped down version of PBR-MW.
Our contribution lies in leveraging this simplicity to develop a specialized method. We identify and implement optimizations unique to radiance surfaces to fully realize the potential of the many-worlds idea.
In radiance surface rendering, image formation is a direct 1:1 mapping between a ray and the nearest intersected surface, while PBR-MW requires a complex nested integration over materials, lighting, and geometry. Our approach to project training images into the scene to establish a radiance field loss depends on this 1:1 mapping and does not efficiently translate to the nested integral structure of a global illumination renderer.
Another important contribution is the introduction of a stochastic background distribution, which enables topological changes and substantially improves reconstruction quality. We show how to cheaply evaluate this strategy in expectation, which is needed to maintain algorithmic parity with NeRF. The associated derivations and simplifications (Appendix A) are specific to radiance surfaces and do not transfer to PBR-MW.
@@ 5.3. Limitations and future work
As shown in Figure 8, our lightweight Laplacian regularization fails when there are insufficient observations to constrain the geometry. Using alternative regularization techniques from state-of-the-art geometry reconstruction methods could help mitigate this issue. Our method also struggles to accurately capture the appearance of conductive materials, which could be addressed by incorporating solutions from prior work [Verbin et al., 2022].
An interesting extension of our work could involve implementing a particle-based storage approach, such as Gaussian splatting [Kerbl et al., 2023]. However, 3D Gaussians are inherently semi-transparent, which conflicts with our assumption of opacity. Future work could explore the use of opaque primitives, such as 2D disks, to replace semi-transparent particles.
@@ 6. Conclusion
The ”many worlds” paradigm—i.e., optimizing a distribution over non-interacting primitives—is relatively new in the field of differentiable rendering. In this paper, we apply it to radiance surface reconstruction, which yields a fast and simple alternative to prior works. Particularly notable is that the derivation began with an evolving surface, yet resulted in remarkably similar equations to volumetric scene reconstructions: ones where losses rather than colors are integrated along rays.
@@
Much engineering has gone into the design of optimized algorithms, regularizers, and heuristics for NeRF-based 3D reconstruction. Our hope is that a large portion of this effort will translate to the radiance field loss and yield state-of-the-art results in the future.
@@ Appendix A Derivation of our loss in NeRF form
Following the introduction of the radiance field loss and the stochastic background, a naive implementation of our method would first sample a surface $\mathcal{M}_{\text{b}}$ from a distribution as the perturbation background, and for each sampled $\mathcal{M}_{\text{b}}$, we need to sample multiple candidates to solve the non-local perturbation problem. Naively applying this strategy would result in inefficient implementation. In the following, we derive an expectation of losses over all potential background surfaces (Figure 13) for a specific candidate.
Let $f_{\mathbf{b}}$ be the probability distribution of the background surface along a ray, where $\int_{0}^{t_{\mathrm{max}}}f_{\mathbf{b}}(t)\,\mathrm{d}t=1$. Without loss of generality, we focus on one candidate position $\mathbf{p}$ at distance $t_{\mathbf{p}}$ along the ray. To avoid clutter, we denote the color error metric $\ell(L,~{}L_{\mathrm{target}})$ as $\ell(L)$. The expectation of all losses in the form of Equation 2 of the main text, local at $\mathbf{p}$, is given by:
@@
This final result is equivalent to the one shown in Figure 2 of the main document. It is now also apparent that we do not need to produce extra samples to evaluate $\mathbb{E}_{t>t_{\mathbf{p}}}[\ell(L_{t})]$.
@@ Appendix B Design space of the background distribution
Section 3.3 of the main document proposes the stochastic background surface. Designing the background surface distribution $f_{\mathbf{b}}$ (Figure 13) involves a tradeoff between exploitation and exploration, offering a wide design space.
@@ Appendix C Additional experiments and results
Applying a Laplacian as a post-processing step requires numerous iterations to address these issues and may degrade geometry in other regions. In contrast, training our algorithm with an exponentially decaying Laplacian is more efficient. Consequently, the results in Figure 12 of the main document are obtained by training our algorithm with an exponentially decaying Laplacian.
@@
Figure 20 shows additional level set rendering results for an outdoor scene. Table 3, Table 4 and Table 5 show the complete PSNR, SSIM and LPIPS results in the Instant NGP codebase. Table 6 shows the complete PSNR results in the ZipNeRF codebase. Table 7 shows the complete Chamfer distance results on the DTU dataset.